### v1

# Interface-induced turbulence in viscous binary fluid mixtures

Abstract
We demonstrate the existence of interface-induced turbulence, an emergent nonequilibrium statistically steady state (NESS) with spatiotemporal chaos, which is induced by interfacial fluctuations in low-Reynolds-number binary-fluid mixtures. We uncover the properties of this NESS via direct numerical simulations (DNSs) of cellular flows in the Cahn-Hilliard-Navier-Stokes (CHNS) equations for binary fluids. We show that, in this NESS, the shell-averaged energy spectrum $E(k)$ is spread over several decades in the wavenumber $k$ and it exhibits a power-law region, indicative of turbulence but without a conventional inertial cascade. To characterize the statistical properties of this turbulence, we compute, in addition to $E(k)$, the time series $e(t)$ of the kinetic energy and its power spectrum, scale-by-scale energy transfer as a function of $k$, and the energy dissipation resulting from interfacial stresses. Furthermore, we analyze the mixing properties of this low-Reynolds-number turbulence via the mean-square displacement (MSD) of Lagrangian tracer particles, for which we demonstrate diffusive behavior at long times, a hallmark of strong mixing in turbulent flows.
I Introduction
Additives can lead to spatiotemporal chaos in a fluid, even when the inertia of the fluid is negligible and the Reynolds number $\mathrm{Re}$ is low. The most notable instance of this is the phenomenon of elastic turbulence in polymer solutions [1, 2, 3]. When elastic polymers are added to a laminar Newtonian solvent, their stretching generates elastic stresses that can trigger instabilities eventually resulting in a chaotic flow, which is characterized by a power-law energy spectrum [1, 4, 5] and strongly intermittent fluctuations [6, 7]. Similar chaotic regimes have been observed in low-inertia wormlike-micellar solutions [8, 9] and in suspensions of microscopic rods [10, 11] and spherical rigid particles [12, 13]. In contrast to conventional hydrodynamic turbulence [14], these examples of low-$\mathrm{Re}$ turbulence do not rely on an energy cascade, through an inertial range, so their main applications are in microfluidics, where additives are employed to enhance mixing [15] as an alternative to passive or active mechanical perturbations [16]. By combining theory and direct numerical simulations (DNSs) we uncover a new type of low-Reynolds turbulence, which is driven by interfacial fluctuations, in viscous binary-fluid mixtures. We call this interface-induced turbulence.
A good understanding of binary-fluid mixtures is crucial for modelling emulsions [17], which have a wide variety of applications in the food [18], cosmetics [19], and pharmaceutical industries [20, 21], often in microfluidic devices, where the enhancement of mixing is of vital importance in many situations. In addition to its practical applications, investigations of low-Re turbulence in systems other than viscoelastic fluids is of fundamental interest in nonlinear physics and fluid dynamics. Therefore, it behooves us to explore the possibility of mixing, induced by low-inertia turbulence, in binary-fluid mixtures. We initiate such an exploration by studying a cellular flow in a two-dimensional (2D) binary-fluid system. The Cahn-Hilliard-Navier-Stokes (CHNS) partial differential equations (PDEs), which couple the fluid velocity $\bm{u}$ with a scalar order parameter $\phi$ that distinguishes between two coexisting phases, provide a natural theoretical framework for such flows. Our investigations, based on direct numerical simulations (DNSs), reveal an emergent nonequilibrium statistically steady state (NESS) with spatiotemporal chaos, which is induced by interfacial fluctuations that destabilize the laminar cellular flow. Thus, we find the elastic-turbulence analog for low-Re binary-fluid mixtures: this leads to a kinetic-energy spectrum $E(k)$, spread over several decades in the wave-number $k$, with a power-law regime that is characterised by an exponent $\simeq-4.5$. By analysing the time dependence of the total kinetic-energy $e(t)$ and its power spectrum, we characterize the transitions from the cellular flow to such turbulence, for which we demonstrate, via a scale-by-scale analysis of the kinetic energy, that there is no significant energy cascade, and therefore the chaotic dynamics is entirely driven by the interfacial stress. Furthermore, we elucidate how such interfacial stress leads to global energy dissipation, even though it is responsibe for both local injection as well as dissipation of energy. Finally, we quantify the mixing properties of interface-induced turbulence by showing that the mean-square-displacement (MSD) of Lagrangian tracers displays long-time diffusive behavior that is similar to its counterpart in inertial turbulence.
II Model
The CHNS PDEs have been used to study multi-fluid flows, which may involve droplet interactions [22, 23, 24, 25, 26], the evolution of antibubbles [27], and phase separation and turbulence in such flows [28, 29, 30, 31]. The two-dimensional incompressible CHNS PDEs are [23, 28, 32]:
$\partial_{t}\phi+\bm{u}\cdot\nabla\phi=M\nabla^{2}\left(\frac{\delta\mathcal{F}}{\delta\phi}\right)\,;$ (1)
$\partial_{t}\omega+\bm{u}\cdot\nabla\omega=\nu\nabla^{2}\omega-\alpha\omega+(\nabla\times\mathcal{\bm{S}}^{\phi})\cdot\hat{e}_{z}+f^{\omega};$ (2)
$\nabla\cdot\bm{u}=0\,;\;\omega=(\nabla\times\bm{u})\cdot\hat{e}_{z}\,;$ (3)
$\nu$, $\alpha$, and $M$ are the kinematic viscosity, friction, and mobility, respectively. We write Eq. (2) in the vorticity-streamfunction ($\omega-\psi$) form, with $\bm{u}=\nabla\times(\psi\hat{e}_{z})$ and $\psi=-\nabla^{-2}\omega$; the surface stress and the Landau-Ginzburg free-energy functional are, respectively,
$\mathcal{\bm{S}}^{\phi}$ $=$ $-\phi\nabla\left(\frac{\delta\mathcal{F}}{\delta\phi}\right)\;\;{\rm{and}}$ (4)
[…]
$\Omega$ is the spatial domain, $\sigma$ is the bare surface tension, and $\epsilon$ the interfacial width. The first term in $\mathcal{F}$ is a double-well potential with minima at $\phi=\pm 1$, which correspond to two bulk phases in equilibrium; the second term is the penalty for interfaces; $\phi$ varies smoothly across an interface.
We study the CHNS PDEs (1)-(5) at low $Re$, with an initially square-crystalline array of vortical structures (a cellular flow), imposed by choosing
$f^{\omega}=\hat{e}_{z}\cdot(\nabla\times\bm{f}^{u})=f_{0}k_{f}[\cos(k_{f}x)+\cos(k_{f}y)]\,,$ (6)
with amplitude $f_{0}$ and wave number $k_{f}$. Such cellular flows have been used to examine the melting of this crystalline array by inertial, elastic, and elasto-inertial turbulence in viscoelastic fluids [33, 34, 35]. For $\alpha=0$ and $\phi(\bm{r})=0$, this system has the stationary solution
$\omega=-\omega_{0}[\cos(k_{f}x)+\cos(k_{f}y)]\,;\;\;\;\omega_{0}=f_{0}/\nu k_{f}\,.$ (7)
The spatiotemporal evolution of this cellular flow depends on the Reynolds, Capillary, Cahn, non-dimensionalised friction, and Péclet numbers that are, respectively,
$\mathrm{Re}$ $=$ $\frac{UL}{\nu}\,,\;\;\mathrm{Ca}=\frac{\nu U}{\sigma}\,,\;\;\mathrm{Cn}=\frac{\epsilon}{L_{0}}\,,$
$\alpha^{\prime}$ $=$ $\alpha T\,,\;\;\mathrm{Pe}=\frac{L^{2}U}{M\sigma},$ (8)
with $U=f_{0}/\nu k_{f}^{2}$, $L=k_{f}^{-1}$, $T=\nu k_{f}/f_{0}$, and $L_{0}$ the side of our square simulation domain. To characterize the mixing because of interface-induced turbulence, we introduce $N_{p}$ tracers into the flow. For tracer $i$ (position $\bm{r}^{i}_{0}$ at time $t_{0}$)
$\frac{d\bm{r}^{i}(t)}{dt}=\bm{v}(\bm{r}^{i},t|\bm{r}^{i}_{0},t_{0})=\bm{u}(\bm{r}^{i},t)\;,$ (9)
where $\bm{r}^{i}(t)$ and $\bm{v}(r^{i},t)$ are the position and velocity of the $i_{th}$ tracer. The mean-squared displacement (MSD) is
$\Delta r^{2}(t)=\langle|\bm{r}(t)-\bm{r}(0)|^{2}\rangle,$ (10)
[…]
Figure 1: Pseudocolor plots: (a) the initial condition for the $\phi$ field [Eq. (11)]; (b) steady-state laminar solution [Eq. (2)] for the vorticity $\omega$, with no droplets or friction; (c) illustrative $\phi$ field in the chaotic regime (Capillary number $\mathrm{Ca}=0.15$); and (d) $\omega$, corresponding to subplot (c), and with the overlaid $\phi=0$ contour lines (in black). See the Supplemental Material [36] for the corresponding movie.
III Numerical methods and initial conditions
We carry out pseudospectral DNSs (parameters in Table I in the Supplemental Material [36]) of the CHNS PDEs (1)-(5), with periodic boundary conditions, a square $(2\pi\times 2\pi)$ box, $512^{2}$ collocation points [37, 23, 38, 39], the $1/2$-dealiasing scheme, and a semi-implicit exponential time difference Runge-Kutta-2 method [40] for time integration. To resolve interfaces, we have three computational grid points in interfacial regions. We obtain $\bm{v}$ from $\bm{u}$ via bilinear interpolation at off-grid points and a first-order Euler scheme for Eq. (9[41, 42]. The initial condition [Fig. 1(a)] comprises $N_{d}$ circular droplets  [^1] We have checked explicitly that our results are independent of the initial arrangements and sizes of the droplets.; droplet $i$, centered at $(x_{i},y_{i})$, has radius $R_{i}$:
$\phi(x,y,t=0)$ $=$ $\sum_{i=1}^{N_{d}}\tanh{\left(\frac{R_{i}-\sqrt{(x-x_{i})^{2}+(y-y_{i})^{2}}}{\epsilon}\right)}\,;$
$\omega(x,y,t=0)$ $=$ $0\,.$ (11)
In Fig. 1(b) we show a pseudocolor plot of $\omega$ for the cellular solution (7), for the single-fluid case ($\alpha=0$).
IV Results
We consider $\mathrm{Re}=1<\mathrm{Re}_{c}=\sqrt{2}$, the single-fluid ($\alpha=0$) critical Reynolds number, given the cellular forcing we use [44]. We choose $\mathrm{Re}<\mathrm{Re}_{c}$ to exclude inertial instabilities, so that we can focus only on interface-induced dynamics. Our DNSs reveal that the second phase leads to interfaces whose fluctuations can destabilise this cellular flow and yield interface-induced turbulence, a NESS with spatiotemporal chaos. In Figs. 1(c) and (d) we present pseudocolor plots, of $\phi$ and $\omega$, respectively, for $\mathrm{Ca}=0.15$, which illustrate the breakdown of the cellular flow in Fig. 1(b) (see also the corresponding movie in the Supplemental Material [36]). Moreover, the time series of the rescaled total energy $e(t)/e_{0}$, with $e_{0}=U^{2}$, shows that, as $\mathrm{Ca}$ is varied, the system undergoes a non-monotonic sequence of transitions between periodic regimes and spatiotemporally chaotic NESSs at low $Re$ (see Fig. 2). In the Supplemental Material [36], we examine the above cellular-to-spatiotemporally chaotic transitions via additional plots of the time series of the total energy $e(t)$, its frequency power spectrum, and pseudocolor plots of the vorticity and the energy spectrum for a wide range of Ca.
$e(t)/e_{0}$
${t/T}$
$\mathrm{Ca}]{\hskip 142.26378pt}$
Figure 2: Enlarged versions of the plots of $e(t)/e_{0}$ versus the scaled time $t/T$: from bottom to top $\mathrm{Ca}=0.1,0.15,0.16,0.18,0.6$. For Ca=0.1, $\mathrm{Ca}=0.1,\,0.16,$ and $\mathrm{Ca}=0.6$, the state shows periodic oscillations in $t$; by contrast, the state is temporally chaotic at $\mathrm{Ca}=0.15$ and $\mathrm{Ca}=0.18$.
Figure 3: Log-log plots of (a) the frequency power spectrum $|\tilde{e}(f)|$ versus the scaled frequency $fT$ for $\mathrm{Ca}=0.15$ and $\mathrm{Ca}=0.16$ and (b) the averaged energy spectrum $E(k)$ in the spatiotemporally chaotic NESS for $\mathrm{Ca}=0.15$; inset: the $k$-dependence of the viscous contribution $2\nu k^{2}E(k)$ and the elastic-transfer term $S(k)$; the black line suggests $E(k)\sim k^{-4.5}$. (c) Plots versus $\mathrm{Ca}$ of $\epsilon_{I}$, $\epsilon_{\nu}$, and $\epsilon_{\phi}$, the contributions of different dissipation terms in Eq. (13); inset: the PDF of the local interfacial-stress contribution $\varepsilon_{\phi}$, for different values of $\mathrm{Ca}$. (d) The pseudocolor plot of $\varepsilon_{\phi}$ for $\mathrm{Ca}=0.15$ at a representative time.
Figure 4: (a) Illustrative plot of a tracer trajectory for $\mathrm{Ca}=0.15$; the colorbar shows the simulation time. (b) Log-log plots of the mean-square displacement (MSD) (averaged over $N_{p}=1024\times 128$ particles), showing short-time ballistic ($\sim t^{2}$) and long-time diffusive (tsimilar-$\sim t$) asymptotes for the spatiotemporally chaotic NESSs ($\mathrm{Ca}=0.15$ and $\mathrm{Ca}=0.18$); only ballistic behavior appears for the periodic time evolution at $\mathrm{Ca}=0.16$.
We turn now to spatiotemporal properties. In Fig. 3(a) we give log-log plots of the power-spectrum of the total energy, $|\tilde{e}(f)|$, versus the normalized frequency $fT$. For $\mathrm{Ca}=0.16$, this spectrum shows a single dominant peak, a signature of temporal periodicity; by contrast, for $\mathrm{Ca}=0.15$, we see a broad power-spectrum, which indicates that $e(t)$ is chaotic. In Fig. 3(b) we characterise the spatial distribution of the kinetic energy via a log-log plot of the shell-averaged energy spectrum $E(k)$ versus the wave-number $k$ (see [36] for the definition), in the spatiotemporally chaotic NESS for $\mathrm{Ca}=0.15$. Over a small range of $k$, $E(k)\sim k^{-4.5}$ [black line in Fig. 3(b)]; this power-law exponent is distinct from the one for 2D fluid turbulence (with an exponent $\simeq-3$ in the forward-cascade regime, if there is no friction [45, 46]). A spectrum steeper than $k^{-3}$ is also a characteristic feature of elastic turbulence in polymer solutions [4, 2]. Unlike inertial fluid turbulence, the low-$Re$ interface-induced turbulence we consider does not show an energy cascade. We demonstrate this via the following scale-by-scale kinetic-energy-budget equation [47, 48, 49]:
$\partial_{t}E(k,t)=T(k,t)-2\nu k^{2}E(k,t)+S(k,t)+F(k,t)\,,$ (12)
where $S(k,t)$ is the contribution of the interfacial stress, $T(k,t)$ is the nonlinear energy transfer, and $F(k,t)$ the energy-injection term (see [36] for the definitions). In the inset of Fig. 3(b), we present the $k$-dependence of the viscous contribution $2\nu k^{2}E(k)$, in blue, and, in red, the contribution of the interfacial stress, $S(k)$; both these terms are equal for all $k$, except at the forcing wave-number $k_{f}=4$. In fluid turbulence, inertia plays a pivotal role in transferring energy from the energy-injection wavenumber(s) to other wavenumbers, and $T(k)$ is non-zero for most $k$. By contrast, in the interface-induced turbulence we consider, inertia is negligible, and energy in wavenumbers other than the injection wavenumber is solely attributable to $S(k)$, which is balanced by $2\nu k^{2}E(k)$; hence, $T(k)$ is negligibly small in Eq. (12). This energy transfer by interfacial stresses is a unique property of low-$Re$ interface-induced turbulence and distinguishes it clearly from fluid turbulence. It is also useful to study the energy-budget equation
$\frac{de(t)}{dt}=\epsilon_{I}-\epsilon_{\nu}-\epsilon_{\phi}\,;$ (13)
$\epsilon_{I}=\langle\bm{f}^{u}\cdot\bm{u}\rangle_{\bm{x}}$ is the mean energy-injection rate, $\epsilon_{\nu}=-\langle\bm{u}\cdot\nu\nabla^{2}\bm{u}\rangle_{\bm{x}}$ is the mean energy-dissipation (viscous) rate, $\epsilon_{\phi}=-\langle\varepsilon_{\phi}\rangle_{\bm{x}}=-\langle\bm{u}\cdot\mathcal{\bm{S}}^{\phi}\rangle_{\bm{x}}$ is the additional mean dissipation because of interfaces, and $\langle\cdot\rangle_{\bm{x}}$ denotes the space average. We plot $\epsilon_{I}$, $\epsilon_{\nu}$, and $\epsilon_{\phi}$ versus $\mathrm{Ca}$ in Fig. 3(c). At intermediate values of $\mathrm{Ca}$, $\epsilon_{\phi}>0$; i.e., globally, the interfacial contribution to the energy budget is dissipative. However, the interfacial stress both injects and dissipates energy locally, as we demonstrate by plotting, in the inset of Fig. 3(c), the probability distribution functions (PDFs) of $\varepsilon_{\phi}$, the local dissipation because of interfaces. The fat tails of this PDF exhibit that $\varepsilon_{\phi}$ shows large fluctuations that are both positive and negative. The pseudocolor plot of $\varepsilon_{\phi}$ for $\mathrm{Ca}=0.15$ in Fig. 3(d) also confirms that $\varepsilon_{\phi}$ is concentrated at the interface between the two fluids. Therefore, the turbulent behavior, which we uncover by the energy-budget analysis (13), is attributable solely to the presence of interfaces in the flow, and is observed at intermediate values of Ca. For low values of $\mathrm{Ca}$ (large $\sigma$), $\epsilon_{\phi}$ is low because the interfaces are so energetic that their energy surpasses the kinetic energy of the flow: thus, droplets coalesce, interfaces do not break-up, and the interfacial length is minimal. For high values of $\mathrm{Ca}$ (low $\sigma$), the interfacial energy is so low that it hardly affects the flow, and the system retains the cellular structure of the applied force; and the energy injection and viscous dissipation balance, i.e., $\epsilon_{I}/\epsilon_{0}=\epsilon_{\nu}/\epsilon_{0}=1$, and $\epsilon_{\phi}/\epsilon_{0}=0$, with $\epsilon_{0}\equiv f_{0}U$.
One of the intriguing properties of interface-induced turbulence is that it enhances mixing even at low $Re$, which makes this phenomenon of great interest for microfluidic applications. We quantify such mixing properties by investigating the dispersion of tracer particles in the flow [Eqs. (9) and (10)]. In Fig. 4(a), we depict a representative tracer trajectory in the spatiotemporally chaotic NESS for $Ca=0.15$; the colorbar shows the simulation time. Initially, the particles get trapped within vortices, but, when an interface moves through these vortices, it facilitates particle transfer to other vortices. We plot the MSD [Eq. (10)] versus $t$ for $\mathrm{Ca}=0.15$, $\mathrm{Ca}=0.16$, and $\mathrm{Ca}=0.18$ in Fig. 4(b). For the chaotic NESSs ($\mathrm{Ca}=0.15$ and $\mathrm{Ca}=0.18$) the small- and large-$t$ asymptotes of the MSD can be fit to the power-law-form $\langle r^{2}(t)\rangle\sim t^{\beta}$, with short-time ballistic behavior $\beta=2$, and long-time diffusive behavior $\beta=1$, because of strong mixing via interface-induced turbulence. If the state is periodic, e.g., for $\mathrm{Ca}=0.16$, the MSD shows only ballistic behavior and then trapping into a vortical cell at longer times.
V Conclusions
We have demonstrated how interfaces in a binary-fluid mixture can disrupt low-$Re$ cellular flows by precipitating instabilities that lead to interface-induced turbulence, the binary-fluid analog of elastic turbulence in fluids with polymer additives [1, 2, 3]. We have explored the transitions from cellular flows, to flows with spatiotemporal crystals, and, eventually, to a NESS with interface-induced turbulence. We have characterised these states via the energy time series $e(t)$, its frequency power spectrum $|\tilde{e}(f)|$, the energy spectrum $E(k)$, the energy budget [Eqs. (12) and 13], and the MSD of Lagrangian tracers [Eqs. (9) and (10)]. The low-$\mathrm{Re}$ interface-induced turbulence that we have uncovered exhibits the following distinctive properties: (a) $|\tilde{e}(f)|$ is significant over a broad range of frequencies $f$; (b) a power-law regime with $E(k)\sim k^{-4.5}$, with a power that is different from its counterpart in 2D fluid turbulence with no friction [45, 46]; (c) a scale-by-scale energy transfer [Eq. (12)] with negligible inertial contribution $T(k)$; (d) an MSD of tracers that crosses over from ballistic to diffusive behaviors, indicating strong mixing. Cellular flows have been used in experimental studies of elastic turbulence [50]; we therefore look forward to experimental confirmations of our predictions for low-$\mathrm{Re}$ interface-induced turbulence in such flows.
Acknowledgements.
We thank K.V. Kiran and S. Mukherjee for valuable discussions. NBP, DV, and RP thank the Indo-French Centre for Applied Mathematics (IFCAM) for support and the Isaac Newton Institute for Mathematical Sciences for support and hospitality during the programme ‘Anti-diffusive dynamics: from sub-cellular to astrophysical scales’ when work on this paper was undertaken (EPSRC grant no EP/R014604/1). This research was supported in part by the International Centre for Theoretical Sciences (ICTS) for the online program - Turbulence: Problems at the Interface of Mathematics and Physics (code: ICTS/TPIMP2020/12). NBP and RP thank the Science and Engineering Research Board (SERB) and the National Supercomputing Mission (NSM), India for support, and the Supercomputer Education and Research Centre (IISc) for computational resources. We thank NEC, India for trial use of the SX-Aurora TSUBASA computer, on which we carried out preliminary DNSs for the 2D CHNS system.
$\mathrm{Ca}]{\hskip 455.24408pt}$
0.010.120.170.190.2(a)(f)(k)(p)(b)(g)(l)(q)(c)(h)(m)(r)(d)(i)(n)(s)(e)(j)(o)(t)
Figure 5: Plots for different Capillary numbers Ca increasing from $0.01$ in the bottom row to $0.2$ in the top row: (a)-(e) The scaled total kinetic energy $e(t)/e_{0}$ versus time $t$, with $e_{0}=U^{2}$, the energy of laminar flow without the second phase, and $U=f_{0}/(\nu k_{f}^{2})$ [see Eq (8) in the main text]. (f)-(j) Plots of the corresponding power spectra $|\tilde{e}(f)|$ of $e(t)/e_{0}$. (k)-(o) Pseudocolor plots of the vorticity $\omega$ overlaid with the $\phi=0$ contour (black lines) at a representative time. (p)-(t) Pseudocolor plots of the energy spectra $\mathcal{E}(k_{x},k_{y})$ at a representative time (other parameters are listed in Table 1).
Appendix A Main definitions
We define the energy spectrum and the shell-averaged energy spectrum as
$\mathcal{E}(\mathbf{k},t)=\left[\hat{\bm{u}}(\mathbf{k},t)\cdot\hat{\bm{u}}(-\mathbf{k},t)\right]$ (14)
and
$E(k,t)=\frac{1}{2}\displaystyle\sum_{k\leq|\bm{k}^{\prime}|<k+1}\left[\hat{\bm{u}}(\mathbf{k}^{\prime},t)\cdot\hat{\bm{u}}(-\mathbf{k}^{\prime},t)\right]\,.$ (15)
In the kinetic-energy-budget equation, the transfer term, the contribution of the interfacial stress, and the energy injection term in the kinetic-energy budget equation are
$T(k,t)\equiv-\mathfrak{R}\left[\displaystyle\sum_{k\leq|\bm{k}^{\prime}|<k+1}[\hat{\bm{u}}(-\bm{k}^{\prime},t)\cdot\bm{P}(\bm{k}^{\prime})\cdot\widehat{(\bm{u}\cdot\nabla\bm{u})}(\bm{k}^{\prime},t)]\right],$ (16)
$S(k,t)\equiv\mathfrak{R}\left[\displaystyle\sum_{k\leq|\bm{k}^{\prime}|<k+1}[\hat{\bm{u}}(-\bm{k}^{\prime},t)\cdot\bm{P}(\bm{k}^{\prime})\cdot\widehat{\mathcal{\bm{S}}^{\phi}}(\bm{k}^{\prime},t)]\right]\,,\;\;$ (17)
and
$F(k,t)\equiv\mathfrak{R}\left[\displaystyle\sum_{k\leq|\bm{k}^{\prime}|<k+1}[\hat{\bm{u}}(-\bm{k}^{\prime},t)\cdot\widehat{f^{u}}(\bm{k}^{\prime},t)]\right]\,.$ (18)
$\mathrm{Ca}]{\hskip 156.49014pt}$
0.160.18(a)(b)(c)(d)(e)(f)(g)(h)
Figure 6: Plots for two illustrative Capillary numbers $\mathrm{Ca}=0.16$ (bottom row) and $\mathrm{Ca}=0.18$ (top row): (a),(e) The scaled total kinetic energy $e(t)/e_{0}$ versus time $t$, with $e_{0}=U^{2}$; (b),(f) power spectra $|\tilde{e}(f)|$ of $e(t)/e_{0}$; (c),(g) pseudocolor plots of the vorticity $\omega$ overlaid with the $\phi=0$ contour (black lines) at a representative time; (d),(h) pseudocolor plots of the 2D energy spectra $\mathcal{E}(k_{x},k_{y})$ at a representative time.
Appendix B Simulation parameters and supplementary analysis
The parameters of the simulation are given in Table 1. Plots of the spatiotemporal properties of the flow are shown in Figs. 5 and 6 for different values of the Capillary number Ca.
    R1     R2     R3     R4     R5     R6     R7     R8     R9     R10
      $\mathrm{Ca}$ 0.01 0.1 0.12 0.15 0.16 0.17 0.18 0.19 0.2 0.6
      $\mathrm{Pe}$ 4 36 42 51 54 58 62 66 72 217
   Nature of state STPOG STPOG STPOG STC STPO STPO STC STC STC STPO
Table 1: The non-dimensional Capillary $\mathrm{Ca}$ and Péclet $\mathrm{Pe}$ numbers for our DNS runs R1-R10 ($\mathrm{Re}=1$, $\mathrm{Cn}=0.006$). We fix the following parameters: $f_{0}=0.1024,\,k_{f}=4,\,\nu=0.04,\,\epsilon=0.03,\,M=6\times 10^{-4},\,$ and $\alpha=0$. The natures of the states [nonequilibrium statistically steady states (NESSs) in the chaotic cases] are given in the last row. STPO: spatially and temporally periodic (i.e., a spatiotemporal crystal); STPOG: STPO but with a grain boundary (G) separating two crystalline parts; STC: Spatiotemporal chaos; TPO: temporal periodic oscillations.
Appendix C Video
  • The video showing the spatiotemporal evolution that corresponds to the pseudocolor plots in Fig. 1 is available at: https://youtu.be/yg28KMcSZQw